OpenStudy (brittany):

Choose the equation of the line that contains the points (8, 0) and (0, 2). y= -1/4x+2,y= -1/4x-2,y= -1/4+2x

OpenStudy (anonymous):

its y=-(1/4)x+2 i hope

OpenStudy (aroub):

2-0/0-8 2/-8 1/-4 y=mx+b , (0,2), m=-1/4 2=-1/4(0)+b 2=b so the equation is y=-1/4x+2

OpenStudy (aroub):

so its the first one! :)

OpenStudy (brittany):

thankyou :)

OpenStudy (aroub):

no problem! :)

OpenStudy (anonymous):

appretiations accepted......................and with a smile